OpenStudy (anonymous):

A balloon is being filled with helium at a rate of 4 ft^3/min. The rate, in square feet per minute, at which the surface are is increasing when the volume is 32(pi)/ 3 is...

OpenStudy (anonymous):

you accidently the problem

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

is it a spherical balloon?

OpenStudy (anonymous):

yeah and i know that the Volume of a sphere is \[4/3 \pi r^2 \]

OpenStudy (anonymous):

A = 4*Pi*r^2

OpenStudy (anonymous):

so thats the surface area...

OpenStudy (anonymous):

volume is (4/3)pi r ^3 but that's not what the question is asking

OpenStudy (anonymous):

oh hold on a sec

OpenStudy (anonymous):

dA/dt = dV/dt * dA/dV

OpenStudy (anonymous):

Volume = Area * r/3

OpenStudy (anonymous):

Area = 3*Volume/r

OpenStudy (anonymous):

r being rate or radius?

OpenStudy (anonymous):

radius

OpenStudy (anonymous):

okay... well i don't know the radius..

OpenStudy (anonymous):

need to replace that as well

OpenStudy (anonymous):

oh okay wait, i know how i can find it

OpenStudy (anonymous):

volume = (4/3)pi r ^3 r = (3/(4Pi) * V)^(1/3)

OpenStudy (anonymous):

Area = 3*Volume/(3/(4Pi) * V)^(1/3)

OpenStudy (anonymous):

area = 3*Volume^(2/3)/(3/(4Pi)^(1/3)

OpenStudy (anonymous):

cancelled volume from the denominator

OpenStudy (anonymous):

now find the derivative to get dA/dV

OpenStudy (anonymous):

to clean it up I'll combine all the numbers

OpenStudy (anonymous):

3/(3/(4Pi))^(1/3)=6^(2/3)*Pi^(1/3)

OpenStudy (anonymous):

area = 6^(2/3)*Pi^(1/3)*V^(2/3)

OpenStudy (anonymous):

dA/dV = (2/3)*6^(2/3)*Pi^(1/3)*V^(-1/3)

OpenStudy (anonymous):

dA/dV = 1 when V= 32*Pi/3

OpenStudy (anonymous):

dA/dt = 4*1 = 4 f^2/min

OpenStudy (anonymous):

well that cleaned up well

OpenStudy (anonymous):

haha, thanks!

OpenStudy (anonymous):

checking to see if i made a mistake anywhere

OpenStudy (anonymous):

seems correct

OpenStudy (anonymous):

you didn't, it was multiple choice and that was an answer.. and the answer key says its right, i just needed to know how to get the answer haha

OpenStudy (anonymous):

that was a difficult question

OpenStudy (anonymous):

yeah, its from my practice AP calc book